OpenStudy (anonymous):

A man is 200 m south of a signal light. He uses a pair of binoculars to observe a car moving at 25 m/s perpendicularly toward the light. At what rate is the man rotating the binoculars at the instant the car is 45 m away from the signal?

OpenStudy (anonymous):

<-45m-> <__CAR O________ /\ | / <----25m/s | | / | | / | | / 200| / | | / | | / | | 0/ Theta \/ |/ MAN

OpenStudy (amistre64):

the tangent of the angle is changing at a certain rate

OpenStudy (anonymous):

What?

OpenStudy (amistre64):

tan(t) = d/200 right?

OpenStudy (amistre64):

d is changing at a rate of 25 m.sec; 200 isnt changing at all ....

OpenStudy (anonymous):

yeah thats right

OpenStudy (amistre64):

tan(theta) = d/200 ; lets derive this with respect to time (d(theta)/dt) sec^2(theta) = (1/200) d(d)/dt ; now we solve for d(theta)/dt

OpenStudy (amistre64):

\[\frac{d(\theta)}{dt}=\frac{25}{200*sec^2(\theta)}\]

OpenStudy (amistre64):

we just need to determine the value of theta at the time indicated and fill it into our equation

OpenStudy (amistre64):

tan(theta) = 45/200 theta = tan-1(45/200)

OpenStudy (amistre64):

another method would be just to determine the value of sec(theta) and square it

OpenStudy (amistre64):

the calculator tells me sec^2(theta) = 1681/1600 so we calculate: 25/(200*(1681/1600)) = 200/1681

OpenStudy (amistre64):

or abt .11898

OpenStudy (amistre64):

we got any options to judge by?

OpenStudy (anonymous):

AMAZING!!!

OpenStudy (anonymous):

that was right! (:

OpenStudy (amistre64):

yay!! :)

OpenStudy (anonymous):

:D